# SFT Doesn’t Always Hurt General Capabilities: Revisiting Domain-Specific Fine-Tuning in LLMs

### v2

@@ C.1.1 MedCalc
We use the MedCalc dataset (khandekar2024medcalc) for medical reasoning tasks. The benchmark provides human-annotated chain-of-thought (CoT) rationales, which we include during training so that the model learns to reason through intermediate steps before producing the final answer. The prompt can be found in Table 6.
@@ C.1.2 ESCI
We use the ESCI dataset (reddy2022shopping) for a multi-class product classification task, where each query–product pair is labeled as Exact, Substitute, Complement, or Irrelevant. From the original dataset, we randomly sample a 5050K subset from the training split and a 1010K subset from the test split. From the training subset, we further hold out 11K examples as a validation set.
@@
The ESCI dataset is highly imbalanced, with the majority of samples belonging to the Exact category (Table 3). This imbalance motivates our choice of balanced accuracy (BACC) as the primary evaluation metric, following prior work on imbalanced classification (xu2024whole; xu2025pisces).
@@ C.1.3 MetaMathQA
MetaMathQA (yu2024metamath) is a large-scale mathematical reasoning dataset containing 395k training examples. Following sanyal2025upweighting, we use MetaMathQA for training and take GSM8K as the target-domain evaluation benchmark. This setup allows us to validate whether our findings hold under large-scale data conditions.
@@ C.1.4 General-Purpose Benchmarks
For the general-purpose benchmarks, we fully follow the default settings and evaluation metrics implemented in the lm-evaluation-harness framework (eval-harness). This ensures consistency with prior work (lin2025rec; sanyal2025upweighting; bansal2025context) and allows for fair comparison of results across different models and training configurations.
@@ C.2 Implementation Details
We conduct all experiments on 16–32 NVIDIA A100 GPUs with 80GB memory. Except for differences in learning rate and loss computation, all experiments share the same training configuration. We adopt the AdamW optimizer (loshchilov2018decoupled) with hyperparameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$, together with a cosine annealing learning-rate schedule. The attention mechanism is implemented using FlashAttention-2 (dao2024flashattention). We set the batch size to 16 for MedCalc and ESCI, and 128 for MetaMathQA. The number of training epochs is 20 for MedCalc and ESCI, and 2 for MetaMathQA. The maximum sequence length is 8192 tokens.
@@ C.4 Effect of KL Regularization
Figure 4: Effect of KL regularization on domain-specific SFT. We follow DeepSeek-R1 (guo2025deepseek) and apply the $k3$ approximation for KL regularization. Results are shown for three learning rates: (a) $1\times 10^{-6}$, (b) $5\times 10^{-6}$, and (c) $2\times 10^{-5}$. Across all settings, KL regularization yields performance that is very close to standard SFT, suggesting limited additional benefit in mitigating general-performance degradation.
We further investigate the effect of KL regularization, a technique recently adopted in DeepSeek-R1 (guo2025deepseek), where a $k3$ approximation is used to estimate the KL term. Following prior work on KL-constrained training (jin2025search; jiang2025deepretrieval; lin2025rec; lin2025training), we add a KL penalty term with coefficient $0.001$ during SFT.
@@ C.6 Performance evolution across training epochs
To better understand how learning rate influences the interaction between domain performance and general performance over the training process, we plot figures of performance vs. training epochs in Figure 6. Figure 6a and 6b report results for Gemma3-4B on the MedCalc benchmark. We make two observations. First, consistent with our main findings, the smallest learning rate achieves strong domain performance while substantially mitigating forgetting on general benchmarks. Second, for each learning rate, the domain performance typically peaks at relatively late epochs. For example, for Gemma3-4B with a learning rate of $1\mathrm{e}{-6}$, the best MedCalc score is reached around epoch 12. This indicates that, during domain-specific SFT, running training for a longer period can continue to improve domain accuracy.
We further examine a larger scale setting with DeepSeek-Coder-7B on MetaMathQA, as shown in Figure 6c and 6d. We observe similar behavior. Domain performance improves steadily and often reaches its peak only after many optimization steps, and smaller learning rates help preserve general capabilities and achieve comparable or even better domain performance. Note that in this case the horizontal axis covers only the first two epochs, but since MetaMathQA is a large scale corpus, even one epoch already corresponds to a large number of parameter update steps. These results confirm that our conclusions are not restricted to small datasets and that the dynamics we describe persist in a large scale training datasets.
@@ Appendix D Details of Token-Adaptive Loss Reweighting
Output: Updated parameters $\theta$
1ex
foreach training step do
   Token NLLs: $\ell_{t}\leftarrow-\log p_{t}$;
 Adaptive weights with lower-bound clipping:
Let $N$ be the number of supervised tokens in the batch;
 Mean (averaged) reweighted loss:
$\mathcal{L}_{\text{TALR}}\;=\;\frac{1}{N}\sum_{t=1}^{N}w_{t}\,(-\log p_{t})$
Parameter update:
$\theta\leftarrow\theta\;-\;\eta\,\nabla_{\theta}\mathcal{L}_{\text{TALR}}$
end foreach
@@ D.3 More Discussions
Comparison with FLOW. It is worth contrasting TALR with FLOW (sanyal2025upweighting), which also reweights losses but in a different manner. First, FLOW operates at the sequence level, whereas TALR works at the token level. Second, FLOW computes static weights only once before training, while TALR dynamically updates weights at every batch with negligible additional cost. As shown in Table 1 and Table 2, TALR consistently outperforms FLOW, which aligns with our expectations. Sequence-level loss can be misleading: for example, even when the overall average sequence loss is small, there may exist a few particularly hard tokens with large losses that are overlooked at the sequence granularity. TALR directly addresses this by reweighting at the token level. Moreover, token difficulty is not fixed—its relative hardness evolves as training progresses, as discussed in Section 4.3. This makes dynamic weighting naturally more advantageous than static approaches.
@@ Appendix E Additional Definitions, Theorems and Proof
IFEval (zhou2023instruction) inst_level_strict_acc, none 0-shot Instruction-following
@@
GSM8K (cobbe2021training) exact_match, flexible-extract 5-shot Mathematical reasoning
@@
HumanEval (chen2021evaluating) pass@1, create_test 0-shot Code generation
@@
HellaSwag (zellers2019hellaswag) acc, none 0-shot Commonsense reasoning
ARC-Easy (Clark2018ThinkYH) acc, none 0-shot Science reasoning
ARC-Challenge (Clark2018ThinkYH) acc, none 0-shot Science reasoning
PIQA (bisk2020piqa) acc, none 0-shot Physical commonsense reasoning
@@
MMLU (hendrycks2020measuring) acc, none 0-shot Multi-domain knowledge understanding
@@
MedCalc (khandekar2024medcalc) Accuracy Medical mathematical reasoning
ESCI (reddy2022shopping) Balanced Accuracy E-commerce product classification

### v3

@@ C.1.1 MedCalc
We use the MedCalc dataset (Khandekar et al., 2024) for medical reasoning tasks. The benchmark provides human-annotated chain-of-thought (CoT) rationales, which we include during training so that the model learns to reason through intermediate steps before producing the final answer. The prompt can be found in Table 6.
@@ C.1.2 ESCI
We use the ESCI dataset (Reddy et al., 2022) for a multi-class product classification task, where each query–product pair is labeled as Exact, Substitute, Complement, or Irrelevant. From the original dataset, we randomly sample a 5050K subset from the training split and a 1010K subset from the test split. From the training subset, we further hold out 11K examples as a validation set.
@@
The ESCI dataset is highly imbalanced, with the majority of samples belonging to the Exact category (Table 3). This imbalance motivates our choice of balanced accuracy (BACC) as the primary evaluation metric, following prior work on imbalanced classification (Xu et al., 2024; 2025).
@@ C.1.3 MetaMathQA
MetaMathQA (Yu et al., 2024) is a large-scale mathematical reasoning dataset containing 395k training examples. Following Sanyal et al. (2025), we use MetaMathQA for training and take GSM8K as the target-domain evaluation benchmark. This setup allows us to validate whether our findings hold under large-scale data conditions.
@@ C.1.4 General-Purpose Benchmarks
For the general-purpose benchmarks, we fully follow the default settings and evaluation metrics implemented in the lm-evaluation-harness framework (Gao et al., 2024). This ensures consistency with prior work (Lin et al., 2025a; Sanyal et al., 2025; Bansal and Sanghavi, 2025) and allows for fair comparison of results across different models and training configurations.
@@ C.2 Implementation Details
We conduct all experiments on 16–32 NVIDIA A100 GPUs with 80GB memory. Except for differences in learning rate and loss computation, all experiments share the same training configuration. We adopt the AdamW optimizer (Loshchilov and Hutter, 2019) with hyperparameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$, together with a cosine annealing learning-rate schedule. The attention mechanism is implemented using FlashAttention-2 (Dao, 2024). We set the batch size to 16 for MedCalc and ESCI, and 128 for MetaMathQA. The number of training epochs is 20 for MedCalc and ESCI, and 2 for MetaMathQA. The maximum sequence length is 8192 tokens.
@@ C.4 Effect of KL Regularization
Figure 5: Effect of KL regularization on domain-specific SFT. We follow DeepSeek-R1 (Guo et al., 2025) and apply the $k3$ approximation for KL regularization. Results are shown for three learning rates: (a) $1\times 10^{-6}$, (b) $5\times 10^{-6}$, and (c) $2\times 10^{-5}$. Across all settings, KL regularization yields performance that is very close to standard SFT, suggesting limited additional benefit in mitigating general-performance degradation.
We further investigate the effect of KL regularization, a technique recently adopted in DeepSeek-R1 (Guo et al., 2025), where a $k3$ approximation is used to estimate the KL term. Following prior work on KL-constrained training (Jin et al., 2025; Jiang et al., 2025; Lin et al., 2025a; b), we add a KL penalty term with coefficient $0.001$ during SFT.
@@ C.6 Performance evolution across training epochs
To better understand how learning rate influences the interaction between domain performance and general performance over the training process, we plot figures of performance vs. training epochs in Figure 7. Figure 7a and 7b report results for Gemma3-4B on the MedCalc benchmark. We make two observations. First, consistent with our main findings, the smallest learning rate achieves strong domain performance while substantially mitigating forgetting on general benchmarks. Second, for each learning rate, the domain performance typically peaks at relatively late epochs. For example, for Gemma3-4B with a learning rate of $1\mathrm{e}{-6}$, the best MedCalc score is reached around epoch 12. This indicates that, during domain-specific SFT, running training for a longer period can continue to improve domain accuracy.
We further examine a larger scale setting with DeepSeek-Coder-7B on MetaMathQA, as shown in Figure 7c and 7d. We observe similar behavior. Domain performance improves steadily and often reaches its peak only after many optimization steps, and smaller learning rates help preserve general capabilities and achieve comparable or even better domain performance. Note that in this case the horizontal axis covers only the first two epochs, but since MetaMathQA is a large scale corpus, even one epoch already corresponds to a large number of parameter update steps. These results confirm that our conclusions are not restricted to small datasets and that the dynamics we describe persist in a large scale training datasets.
@@ D.3 More Discussions
Comparison with FLOW. It is worth contrasting TALR with FLOW (Sanyal et al., 2025), which also reweights losses but in a different manner. First, FLOW operates at the sequence level, whereas TALR works at the token level. Second, FLOW computes static weights only once before training, while TALR dynamically updates weights at every batch with negligible additional cost. As shown in Table 1 and Table 2, TALR consistently outperforms FLOW, which aligns with our expectations. Sequence-level loss can be misleading: for example, even when the overall average sequence loss is small, there may exist a few particularly hard tokens with large losses that are overlooked at the sequence granularity. TALR directly addresses this by reweighting at the token level. Moreover, token difficulty is not fixed—its relative hardness evolves as training progresses, as discussed in Section 4.3. This makes dynamic weighting naturally more advantageous than static approaches.
@@ Appendix E Additional Definitions, Theorems and Proof
IFEval (Zhou et al., 2023) inst_level_strict_acc, none 0-shot Instruction-following
@@
GSM8K (Cobbe et al., 2021) exact_match, flexible-extract 5-shot Mathematical reasoning
@@
HumanEval (Chen et al., 2021) pass@1, create_test 0-shot Code generation
@@
HellaSwag (Zellers et al., 2019) acc, none 0-shot Commonsense reasoning
ARC-Easy (Clark et al., 2018) acc, none 0-shot Science reasoning
ARC-Challenge (Clark et al., 2018) acc, none 0-shot Science reasoning
PIQA (Bisk et al., 2020) acc, none 0-shot Physical commonsense reasoning
@@
MMLU (Hendrycks et al., 2020) acc, none 0-shot Multi-domain knowledge understanding
@@
MedCalc (Khandekar et al., 2024) Accuracy Medical mathematical reasoning
ESCI (Reddy et al., 2022) Balanced Accuracy E-commerce product classification